Given the following information. Percent of capital structure: Debt 10 % Preferred stock 5 Common equity 85 Additional information: Bond coupon rate 13 % Bond yield 11 % Dividend, expected common $7.00 Dividend, preferred $14.00 Price, common $70.00 Price, preferred $110.00 Flotation cost, preferred $2.50 Corporate growth rate 4 % Corporate tax rate 30 % Calculate the weighted average cost of capital for Genex Corporation. Line up the calculations in the order shown in Table 11-1. (Do not round your intermediate calculations and round your final answers to 2 decimal places.) Weighted Cost Debt (Kd) % Preferred stock (Kp) Common equity (Ke) Weighted average cost of capital (Ka) %
Answer #1

Solution:

Cost of debt = Bond yield *(1- tax rate)

Cost of debt = 11%*(1-0.30)= 7.7%

Cost of preferred stock = preferred dividend/( price - flotation cost)

Cost of preferred stock =[ $14 /(110-2.50)]*100= 13.02%

Cost of common equity =( D1/P0)+g

Cost of common equity = (7/70)+0.04= 14%

Weighted average cost of capital = 7.7%*0.1 + 13.02%*0.05 + 14%*0.85= 13.32%